The perimeter of an isosceles triangle is 71 centimeters. The measure of one of the sides is 22 centimeters. What is one possibility of the measures of the other two sides?
Answer by edjones(8007)

There are only 2 possibilities.
22, 22, 71-44=27
.
71-22=49
49/2=24.5
22, 24.5, 24.5
.
Ed